Find the length of each side of the rectangle so that it has the given perimeter.

2x - 5
x
Perimeter = 50 in.


Answer by checkley75(3666)

2x+2(2x-5)=50
2x+4x-10=50
6x=50=10
6x=60
x=10 answer
2*10-5=20-5=15 answer
proof
2*10+2*15=50
20+30=50
50=50